OpenStudy (kpop4life123):

Find the value of the following expression: (2^8 ⋅ 3^–5 ⋅ 6^0)^–2 ⋅ (3^-2 / 2^3)^4 ⋅ 2^28 Write your answer in simplified form. Show all of your steps.

OpenStudy (anonymous):

First turn the exponents into numbers... for example 2^8 = 256, you can use a calculator.

OpenStudy (kpop4life123):

oh ok

OpenStudy (anonymous):

Next multiply/divide (in parenthesis), then add/subtract (in parenthesis), then find the value of those by the exponent out side of the parenthesis

OpenStudy (anonymous):

then just follow PEMDAS

OpenStudy (kpop4life123):

is the answer 9?

OpenStudy (kpop4life123):

is it right?

OpenStudy (anonymous):

let me check. One second!

OpenStudy (anonymous):

Sorry, it may take a bit longer- i tried to do it with decimals xD

OpenStudy (kpop4life123):

haha its ok

OpenStudy (anonymous):

ugh, the 3^-5 is being difficult to work with...

OpenStudy (lolacole12):

(28×3−5)−2 multiply the exponents at get 2−1635

OpenStudy (kpop4life123):

i got 0.0041 for 3^-5

OpenStudy (anonymous):

so you're rounding?

OpenStudy (anonymous):

i made it \[\frac{ 256 }{ 243 }\] instead

OpenStudy (anonymous):

(including multiplying by 256)

OpenStudy (kpop4life123):

OMG this problem is hard to do!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yes, very hard...

OpenStudy (anonymous):

im willing to try tho =D

OpenStudy (kpop4life123):

look what i got for 2^28 = 268435456

OpenStudy (kpop4life123):

thank you so much btw!

OpenStudy (anonymous):

Yes, I got the same- Right now im at (256/243)^-2*(1/26873856) *268435456

OpenStudy (anonymous):

No problem! I like to help :)

OpenStudy (kpop4life123):

:)

OpenStudy (anonymous):

Now im at 0.00001693508*(1/26873856) *268435456

OpenStudy (kpop4life123):

omg im on my 4th try ive been doing this problem for a long time

OpenStudy (anonymous):

okay, now its 0.00001693508*9.98872123152

OpenStudy (kpop4life123):

so this whole thing is an answer ? i thought it was 9

OpenStudy (anonymous):

0.000169159793153 This is my answer, I deleted the last one because I realized it was wrong xD

OpenStudy (anonymous):

I don't know how you got nine, mind explaining?

OpenStudy (kpop4life123):

hahaha i typed in everything in the calculator

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Oh yea... it gave me that too...

OpenStudy (kpop4life123):

i still dont know what to enter in for an explanation

OpenStudy (anonymous):

Ask your teacher man, this is nuts xD

OpenStudy (anonymous):

The calculator only did ^-2 for the 6^0, instead of doing it to the whole equation

OpenStudy (kpop4life123):

ok ill type in everything i did and you just check ok :)

OpenStudy (anonymous):

Oh you know what?... it is nine! xD, or im pretty sure it is... The first part adds up to about .9, which means that if the other part is *10, then it will be 9. The second part is 9.9, but account for rounding, and .9*10=9!

OpenStudy (kpop4life123):

im sooo confused but are you sure its 9

OpenStudy (anonymous):

I'm very convinced... Not smart enough to be sure... my suggestion is you get qualified help

OpenStudy (kpop4life123):

ok thanks for your help :) im going to repost this again to see if anyone knows it or im just going to turn in what i got

OpenStudy (kpop4life123):

but still thanks you so much for helping me :)

OpenStudy (anonymous):

yeah, good idea, your welcome ;)